OpenStudy (anonymous):

What is the equation of a horizontal line passing through the point (2, 10)? y = 12 x = −8 y = 10 x = 2

OpenStudy (anonymous):

@ganeshie8 10?

ganeshie8 (ganeshie8):

10 is a number, not an equation of a line

OpenStudy (anonymous):

y = 10 .-.

ganeshie8 (ganeshie8):

:) Correct !

ganeshie8 (ganeshie8):

for a horizontal line, y coordinate is fixed.

OpenStudy (anonymous):

okay, thanks. i have a couple others i need help on if you wouldn't mind

OpenStudy (anonymous):

Find the average rate of change for the given function from x = 1 to x = 2. -2 \[- \frac{ 1 }{ 2 }\] \[\frac{ 1 }{ 2 }\] 2